# Scalarisation-based risk concepts for robust multi-objective optimisation

### v2

@@ 1 Introduction
Robust multi-objective optimisation is concerned with the problem of optimising a vector-valued function under uncertainty. In this paper, we will consider optimising a bounded objective function $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$, which is defined over a $D$-dimensional input space $\mathbb{X}\subseteq\mathbb{R}^{D}$ and a $W$-dimensional uncertain parameter space $\Xi\subseteq\mathbb{R}^{W}$. Here, the inputs $\mathbf{x}\in\mathbb{X}$ are used to denote all the parameters that can be controlled by the decision maker. Whilst the uncertain parameters $\boldsymbol{\xi}\in\Xi$ are used to denote all the other variables in the problem, which effect the value of the objective function, but are subject to uncertainty or randomness. Notably, these uncertain parameters might not be controllable. As a result, the goal of robust multi-objective optimisation is to identify a collection of controllable inputs whose distribution of output vectors are in some sense robust to the potential variation in these uncertain parameters. That is, we aim to solve the problem
@@
The majority of existing work in robust multi-objective optimisation has largely focussed on defining the robust optimum via a first principles Pareto ordering style approach where one defines a partial ordering over the space of inputs—see for instance the work by Ide and Köbis (2014); Ide and Schöbel (2016) and Botte and Schöbel (2019) for a review. Conceptually, this partial ordering determines whether an input $\mathbf{x}\in\mathbb{X}$ is considered to be more robust than another input $\mathbf{x}^{\prime}\in\mathbb{X}$ with respect to the objective function $f$. The solution to the corresponding robust optimisation problem is then given by the collection of inputs which are not robustly dominated by any other feasible input. Consequently, one is then typically interested in coming up with a principled computational strategy in order to identify these robust points. The most prevalent approach to accomplish this is the scalarisation strategy. In this approach, one recasts the corresponding robust multi-objective optimisation problem into a sensible collection of scalar-valued optimisation problems which can then be solved using traditional techniques in order to obtain an approximation to the robust set of interest. Although theoretically motivated, the first principles approach is rather rigid in practice because whenever we want to use a new partial ordering, we would have to manually identify and derive a new set of results. In this work, we take a more efficient approach to robust multi-objective optimisation and define robustness directly using scalarisation functions. That is, we define the solution of the robust multi-objective optimisation (1) as being equivalent to solving a collection of general robust scalarised problems that can be readily solved using standard scalar-valued optimisation techniques.
@@
There are the two essential ingredients to our robust multi-objective optimisation strategy: scalarisation and robustification. Scalarisation refers to the procedure of mapping a vector-valued objective function to a family of scalar-valued rewards. Whilst robustification is concerned with the process of quantifying the robust value of a set of possible points. In this work, we focus on the use of risk functionals, otherwise known as risk measures, in order to perform the robustification procedure. These risk inspired concepts of robustness are very well-established in the robust single-objective optimisation literature (Ben-Tal et al., 2009; Bertsimas et al., 2011; Rahimian and Mehrotra, 2019), but have not been adequately studied in the multi-objective setting. Our work aims to address this gap by providing a generic methodology for general risk functionals. Similarly, for the scalarisation procedure, our methodology is general and can accommodate for any arbitrary family of scalarisation functions. Nevertheless, as a working example, we will often use the length scalarisation functions. These are a variant of the Chebyshev scalarisation functions that are known to exhibit many intuitive and useful properties which will be discussed throughout (Ishibuchi et al., 2009; Deb and Jain, 2012; Tu et al., 2024b).
@@ 1.1 Main contributions
As part of this work, we have made several novel and insightful contributions to the study of robust multi-objective optimisation. Firstly, we have introduced and formalised two general computationally efficient approaches to robust multi-objective optimisation: the RTS approach (Section 3.1) and the STR approach (Section 3.2). As we motivate in Section 3.3.1, these two approaches are very flexible and they encompass many existing robust multi-objective solution strategies as special cases. Secondly, we present a detailed discussion on when one should use an RTS or STR approach (Section 3.3.2) and when these approaches are equivalent (Section 3.3.3). Thirdly, inspired by the recent results by Tu et al. (2024b), we showcase in Section 4 how it is possible to define a valid Pareto front surface which reflects the robust trade-off that is happening under both the RTS and STR paradigm for the length scalarised problems. Specifically, we derive the novel concept of the RTS front (Definition 4.3) and STR front (Definition 4.4), which are polar surfaces described by the robust length scalarised values. Fourthly, in Section 5, we present a general class of performance metrics that can be used under the RTS and STR setting in order to assess the performance of any robust multi-objective algorithm. In particular, we generalise the construction of the R2 utility (Hansen and Jaszkiewicz, 1998) to the RTS and STR setting.
@@ 2.1 Multi-objective optimisation
The Pareto domination region of a set of vectors $A\subset\mathbb{R}^{M}$ is defined as the collection of vectors which dominates (or is dominated by) this set, that is
@@ 2.2 Scalarisation perspective
The standard strategy that is often used to solve (2) is the scalarisation method. In this approach, we recast the vector-valued optimisation problem of interest into a collection of scalar-valued optimisation problems
@@
where $s_{\boldsymbol{\theta}}:\mathbb{R}^{M}\rightarrow\mathbb{R}$ is a parametric scalarisation function that is indexed by scalarisation parameters $\boldsymbol{\theta}\in\Theta$. Notably, if one is interested in capturing the Pareto front using (3), then this would require us to define a suitable set of scalarisation functions $\{s_{\boldsymbol{\theta}}\}_{\boldsymbol{\theta}\in\Theta}$ which ensures that the resulting solutions are Pareto optimal. For instance, this property is usually assured by restricting our attention to scalarisation functions that are monotonically increasing (Tu et al., 2024a, Definition 2.5). In practice however, a decision maker might not necessarily be interested in just targetting the Pareto front. In particular they might be open to the use of non-monotonic scalarisation functions, which can be used to target desirable points that are not necessarily Pareto optimal. For example, they might consider the use of distance-based scalarisation functions in order to target a set of desired specifications.
@@
As a working example, we will often use the length scalarisation function, which is a specific reparameterisation of the Chebyshev scalarisation function (Ishibuchi et al., 2009; Deb and Jain, 2012; Tu et al., 2024b):
@@
for any vector $\mathbf{y}\in\mathbb{R}^{M}$, where $\boldsymbol{\eta}\in\mathbb{R}^{M}$ denotes a fixed reference vector and $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$ denotes a positive unit vector that lives in the space of all positive unit vectors $\mathcal{S}^{M-1}_{+}:=\{\mathbf{y}\in\mathbb{R}_{>0}^{M}:||\mathbf{y}||_{L^{2}}=1\}$. Conceptually, for any vector $\mathbf{y}$ that strongly dominates the reference vector $\boldsymbol{\eta}$, the length scalarisation function returns the projected length of this vector along the line $\{\boldsymbol{\eta}+t\boldsymbol{\lambda}:t\in\mathbb{R}\}$, that is, $s^{\textnormal{Len}}_{(\boldsymbol{\eta},\boldsymbol{\lambda})}(\mathbf{y})=\max\{t>0:\boldsymbol{\eta}+t\boldsymbol{\lambda}\preceq\mathbf{y}\}$ for any $\mathbf{y}\in\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$—see the work by Tu et al. (2024b) for more details for this construction. We now recall a key result from the multi-objective optimisation literature, which states that all weakly Pareto optimal solutions of (2) can be recovered by solving a set of length scalarised problems (3). For some additional intuition, we have also included Figure 1, which showcases the geometric idea behind this result in two dimensions.
@@ 2.3 Risk functionals
Most existing work in robust multi-objective optimisation has focussed on the simple scenario where each uncertain parameter $\boldsymbol{\xi}\in\Xi$ is given equal weight. In this work, we consider a more general treatment where the uncertain parameter is treated as a random variable that is defined sensibly over a standard probability space. Precisely, we suppose that the uncertain parameter $\boldsymbol{\xi}\in\Xi$ is distributed according to some probability density $p\in\mathcal{P}[\Xi]$ that could potentially be known but might in general be unknown. Equipped with this set-up, the value of an input $\mathbf{x}\in\mathbb{X}$ will no longer be determined by its equally weighted output set (7), but instead will be defined using a sensible weighted alternative. More specifically, we will consider the use of risk functionals (Rüschendorf, 2013, Part 2) in order to obtain these proxy set of values. For this reason, we will now give a brief overview on the concepts of a univariate risk functional (Section 2.3.1) and a multivariate risk functional (Section 2.3.2). In Figure 2, we give an illustration of some univariate and multivariate risk functionals for a simple one-dimensional and two-dimensional example, respectively.
@@ 2.3.1 Univariate risk functionals
In the following sections, we will consider the problem of maximising the output of a risk functional. We will often refer to the resulting output of a univariate risk functional as the risk-adjusted or risk-penalised objective value. Naturally, we want these risk functionals to satisfy some sensible properties in order for the corresponding risk-penalised optimisation problem to make sense from a modelling perspective. For this reason, we focus our attention on partially coherent univariate risk functional (Definition 2.4). By design, these functionals satisfy four additional properties which are sufficient enough to prove many of the results which we state and describe later in Section 4. Note that all of the univariate risk functionals listed above (Examples 2.2, 2.3, 2.4 and 2.5) are partially coherent.
@@
A partially coherent univariate risk functional becomes a coherent univariate risk functional, as defined by Artzner et al. (1999), if it satisfies an additional property known as sub-additivity: $\rho[a+b]\leq\rho[a]+\rho[b]$ for any functions $a:\Xi\rightarrow\mathbb{R}$ and $b:\Xi\rightarrow\mathbb{R}$. By a well-known duality result (Artzner et al., 1999, Proposition 4.1), it is known that any coherent risk univariate functional can be written as a distributionally robust functional (Example 2.5) with some uncertainty set $\mathcal{U}\subseteq\mathcal{P}[\Xi]$.
@@ 2.3.2 Multivariate risk functionals
where the supremum of a set of vectors is defined with respect to the weak Pareto partial ordering on vectors. At the lowermost case, when $\alpha\downarrow 0$, we obtain the weak Pareto front of the domination region of the essential output set: $\mathcal{Y}^{\textnormal{weak}}[\mathbb{D}_{\preceq}[h(\Xi^{\textnormal{ess}})]]$ where $\Xi^{\textnormal{ess}}:=\textnormal{ess supp}(p)$ is the essential support of $p$. Whilst in the uppermost case, when $\alpha\uparrow 1$, we obtain the weak Pareto front of the domination region of the worst-case vector: $\mathcal{Y}^{\textnormal{weak}}[\mathbb{D}_{\preceq}[\boldsymbol{\rho}^{\textnormal{WorstCase}}[h]]]$ where $\mathcal{E}=\Xi^{\textnormal{ess}}$.
@@
Given a collection of bounded univariate risk functionals $\{\rho^{(m)}$, …, $\rho^{(M)}\}$, the component-wise (CW) multivariate risk functional maps any vector-valued function $h:\Xi\rightarrow\mathbb{R}^{M}$ to the singleton set comprised of its corresponding component-wise risk-adjusted vector
@@ 3.1 Robustify then scalarise
We now introduce a general class of robust multi-objective optimisation problems which we refer to as the robust then scalarise problems. A scalar-valued optimisation problem belongs to this class if it can be written in the form
@@ 3.2 Scalarise then robustify
where $s_{\boldsymbol{\theta}}:\mathbb{R}^{M}\rightarrow\mathbb{R}$ denotes a scalarisation function and $\rho$ denotes a univariate risk functional (Section 2.3.1). Conceptually, an STR problem associates each input $\mathbf{x}\in\mathbb{X}$ with its risk-adjusted scalarised value: $\rho[s(f(\mathbf{x},\cdot))]\in\mathbb{R}$. The solution of an STR problem is then given by the subset of inputs which map to the largest possible risk-adjusted scalarised value. We refer to these solutions as the STR robust solutions of $f$ (with respect to $s_{\boldsymbol{\theta}}$ and $\rho$). Similar with the RTS method (Section 3.1), the STR method works by reframing the robust multi-objective optimisation problem (1) into a collection of STR problems. As before, each scalarised problems corresponds to one specific preference $\boldsymbol{\theta}\in\Theta$ and one set of solutions that are STR robust with respect to this preference. By solving a collection of these problems, one hopes to identify a collection of STR robust solutions that are desirable for the decision maker. Moreover, as these problems are all scalar-valued, one can in theory appeal to many standard techniques from single-objective optimisation in order to solve them.

### v4

@@ 1 Introduction
Robust multi-objective optimisation is concerned with the problem of optimising a vector-valued function under uncertainty. In this paper, we will consider optimising a bounded objective function $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$, which is defined over a $D$-dimensional input space $\mathbb{X}\subseteq\mathbb{R}^{D}$ and a $W$-dimensional uncertain parameter space $\Xi\subseteq\mathbb{R}^{W}$. Here, the inputs $\mathbf{x}\in\mathbb{X}$ are used to denote all the parameters that can be controlled by the decision maker. The uncertain parameters $\boldsymbol{\xi}\in\Xi$ are used to denote all the other variables in the problem, which affect the value of the objective function, but are subject to uncertainty or randomness. Notably, these uncertain parameters might not be controllable. As a result, the goal of robust multi-objective optimisation is to identify a collection of controllable inputs whose distribution of output vectors are in some sense robust to the potential variation in these uncertain parameters. That is, we aim to solve the problem
@@
The majority of existing work in robust multi-objective optimisation has largely focussed on defining the robust optimum via a first principles Pareto ordering style approach where one defines a partial ordering over the space of inputs—see for instance the work by Ide and Köbis (2014); Ide and Schöbel (2016) and Botte and Schöbel (2019) for a review. Conceptually, this partial ordering determines whether an input $\mathbf{x}\in\mathbb{X}$ is considered to be more robust than another input $\mathbf{x}^{\prime}\in\mathbb{X}$ with respect to the objective function $f$. The solution to the corresponding robust optimisation problem is then given by the collection of inputs which are not robustly dominated by any other feasible input. Consequently, one is then typically interested in coming up with a principled computational strategy in order to identify these robust points. The most prevalent approach to accomplish this is the scalarisation strategy. In this approach, one recasts the corresponding robust multi-objective optimisation problem into a sensible collection of scalar-valued optimisation problems; these problems can then be solved using traditional techniques in order to obtain an approximation to the robust set of interest. Although theoretically motivated, the first principles approach is rather rigid in practice because whenever we want to use a new partial ordering, we would have to manually identify and derive a new set of results. In this work, we take a more efficient approach to robust multi-objective optimisation and define robustness directly using scalarisation functions. That is, we define the solution of the robust multi-objective optimisation (1) as being equivalent to solving a collection of general robust scalarised problems that can be readily solved using standard scalar-valued optimisation techniques.
@@
There are the two essential ingredients to our robust multi-objective optimisation strategy: scalarisation and robustification. Scalarisation refers to the procedure of mapping a vector-valued objective function to a family of scalar-valued rewards. Meanwhile, robustification is concerned with the process of quantifying the robust value of a set of possible points. In this work, we focus on the use of risk functionals, otherwise known as risk measures, in order to perform the robustification procedure. These risk-inspired concepts of robustness are very well established in the robust single-objective optimisation literature (Quaranta and Zaffaroni, 2008; Ben-Tal et al., 2009; Bertsimas et al., 2011; Rahimian and Mehrotra, 2019), but have not been adequately studied in the multi-objective setting. Our work aims to address this gap by providing a generic methodology for general risk functionals. Similarly, for the scalarisation procedure, our methodology is general and can accommodate for any arbitrary family of scalarisation functions such as the Gerstewitz functional (Gerstewitz, 1983; Tammer and Weidner, 2020). Nevertheless, as a working example, we will often use the length scalarisation functions. These are a variant of the Chebyshev scalarisation functions that are known to exhibit many intuitive and useful properties which will be discussed throughout (Ishibuchi et al., 2009; Deb and Jain, 2012; Tu et al., 2024).
@@ 1.1 Main contributions
As part of this work, we have made several novel and insightful contributions to the study of robust multi-objective optimisation. Firstly, we have introduced and formalised two general computationally efficient approaches to robust multi-objective optimisation: the RTS approach (Section 3.1) and the STR approach (Section 3.2). As we motivate in Section 3.3.1, these two approaches are very flexible and they encompass many existing robust multi-objective solution strategies as special cases. Secondly, we present a detailed discussion on when one should use an RTS or STR approach (Section 3.3.2) and when these approaches are equivalent (Section 3.3.3). Thirdly, inspired by the recent results by Tu et al. (2024), we showcase in Section 4 how it is possible to define a valid Pareto front surface which reflects the robust trade-off that is happening under both the RTS and STR paradigm for the length scalarised problems. Specifically, we derive the novel concept of the RTS front (Definition 4.3) and STR front (Definition 4.4), which are polar surfaces described by the robust length scalarised values. Fourthly, in Section 5, we present a general class of performance metrics that can be used under the RTS and STR setting in order to assess the performance of any robust multi-objective algorithm. In particular, we generalise the construction of the R2 utility (Hansen and Jaszkiewicz, 1998) to the RTS and STR setting.
@@ 2.1 Multi-objective optimisation
The Pareto domination (or dominated) region of a set of vectors $A\subseteq\mathbb{R}^{M}$ is defined as the collection of vectors which dominates (or is dominated by) this set, that is
@@ 2.2 Scalarisation perspective
The standard strategy that is often used to solve (2) is a scalarisation method. In this approach, we recast the vector-valued optimisation problem of interest into a collection of scalar-valued optimisation problems
@@
where $s_{\boldsymbol{\theta}}:\mathbb{R}^{M}\rightarrow\mathbb{R}$ is a parametric scalarisation function that is indexed by scalarisation parameters $\boldsymbol{\theta}\in\Theta$. Notably, if one is interested in capturing the Pareto front using (3), then this would require us to define a suitable [^1] In theory, one might desire a family of scalarisation functions that can target all Pareto optimal solutions. In practice however, one can usually only recover a finite approximation to the optimal set and therefore this requirement of complete coverage might not be practically necessary. set of scalarisation functions $\{s_{\boldsymbol{\theta}}\}_{\boldsymbol{\theta}\in\Theta}$ which ensures that the resulting solutions are Pareto optimal. For instance, this former property is usually assured by restricting our attention to scalarisation functions that are monotonically increasing (Tu et al., 2025, Definition 2.5). In practice however, a decision maker might not necessarily be interested in just targetting the Pareto front. In particular they might be open to the use of non-monotonic scalarisation functions, which can be used to target desirable points that are not necessarily Pareto optimal. For example, they might consider the use of distance-based scalarisation functions in order to target a set of desired specifications.
@@
As a working example, we will often use the length scalarisation function, which is a specific reparameterisation of the Chebyshev scalarisation function (Ishibuchi et al., 2009; Deb and Jain, 2012; Tu et al., 2024):
@@
for any vector $\mathbf{y}\in\mathbb{R}^{M}$, where $\boldsymbol{\eta}\in\mathbb{R}^{M}$ denotes a fixed reference vector and $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$ denotes a positive unit vector that lives in the space of all positive unit vectors $\mathcal{S}^{M-1}_{+}:=\{\mathbf{y}\in\mathbb{R}_{>0}^{M}:||\mathbf{y}||_{L^{2}}=1\}$. Conceptually, for any vector $\mathbf{y}$ that strongly dominates the reference vector $\boldsymbol{\eta}$, the length scalarisation function returns the projected length of this vector along the line $\{\boldsymbol{\eta}+t\boldsymbol{\lambda}:t\in\mathbb{R}\}$, that is, $s^{\textnormal{Len}}_{(\boldsymbol{\eta},\boldsymbol{\lambda})}(\mathbf{y})=\max\{t>0:\boldsymbol{\eta}+t\boldsymbol{\lambda}\preceq\mathbf{y}\}$ for any $\mathbf{y}\in\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$—see the work by Tu et al. (2024) for more details for this construction. We now recall a key result from the multi-objective optimisation literature, which states that all weakly Pareto optimal solutions of (2) can be recovered by solving a set of length scalarised problems (3). For some additional intuition, we have also included Figure 1, which showcases the geometric idea behind this result in two dimensions.
@@ 2.3 Risk functionals
Most existing work in robust multi-objective optimisation has focussed on the simple scenario where each uncertain parameter $\boldsymbol{\xi}\in\Xi$ is given equal weight. In this work, we consider a more general treatment where the uncertain parameter is treated as a random variable that is defined sensibly over a standard probability space. More precisely, we suppose that the uncertain parameter $\boldsymbol{\xi}\in\Xi$ is distributed according to some probability density $p\in\mathcal{P}[\Xi]$ that could potentially be known but might in general be unknown. Equipped with this set-up, the value of an input $\mathbf{x}\in\mathbb{X}$ will no longer be determined by its equally weighted output set (7), but instead will be defined using a sensible weighted alternative. More specifically, we will consider the use of risk functionals (Rüschendorf, 2013, Part 2) in order to obtain these proxy set of values. For this reason, we will now give a brief overview on the concepts of a univariate risk functional (Section 2.3.1) and a multivariate risk functional (Section 2.3.2). In Figure 2, we give an illustration of some univariate and multivariate risk functionals for a simple one-dimensional and two-dimensional example, respectively.
@@ 2.3.1 Univariate risk functionals
In the following sections, we will consider the problem of maximising the output of a risk functional. We will often refer to the resulting output of a univariate risk functional as the risk-adjusted or risk-penalised objective value. Naturally, we want these risk functionals to satisfy some sensible properties in order for the corresponding risk-penalised optimisation problem to make sense from a modelling perspective. For this reason, we focus our attention on partially coherent univariate risk functional (Definition 2.4). By design, these functionals satisfy four additional properties which are sufficient to prove many of the results which we state and describe later in Section 4. Note that all of the univariate risk functionals listed above (Examples 2.2, 2.3, 2.4 and 2.5) are partially coherent.
@@
A partially coherent univariate risk functional becomes a coherent univariate risk functional, as defined by Artzner et al. (1999), if it satisfies an additional property known as sub-additivity: $\rho[a+b]\leq\rho[a]+\rho[b]$ for any functions $a:\Xi\rightarrow\mathbb{R}$ and $b:\Xi\rightarrow\mathbb{R}$. By a well-known duality result (Artzner et al., 1999, Proposition 4.1), it is known that any coherent univariate risk functional can be written as a distributionally robust functional (Example 2.5) with some uncertainty set $\mathcal{U}\subseteq\mathcal{P}[\Xi]$.
@@ 2.3.2 Multivariate risk functionals
where the supremum of a set of vectors is defined with respect to the weak Pareto partial ordering on vectors. At the lowermost case, when $\alpha\downarrow 0$, we obtain the weak Pareto front of the domination region of the essential output set: $\mathcal{Y}^{\textnormal{weak}}[\mathbb{D}_{\preceq}[h(\Xi^{\textnormal{ess}})]]$ where $\Xi^{\textnormal{ess}}:=\textnormal{ess supp}(p)$ is the essential support of $p$. In the uppermost case, when $\alpha\uparrow 1$, we obtain the weak Pareto front of the domination region of the worst-case vector: $\mathcal{Y}^{\textnormal{weak}}[\mathbb{D}_{\preceq}[\boldsymbol{\rho}^{\textnormal{WorstCase}}[h]]]$ where $\mathcal{E}=\Xi^{\textnormal{ess}}$.
@@
Given a collection of bounded univariate risk functionals $\{\rho^{(1)}$, …, $\rho^{(M)}\}$, the component-wise (CW) multivariate risk functional maps any vector-valued function $h:\Xi\rightarrow\mathbb{R}^{M}$ to the singleton set comprised of its corresponding component-wise risk-adjusted vector
@@ 3.1 Robustify then scalarise
We now introduce a general class of robust multi-objective optimisation problems which we refer to as the robustify then scalarise problems. A scalar-valued optimisation problem belongs to this class if it can be written in the form
@@ 3.2 Scalarise then robustify
where $s_{\boldsymbol{\theta}}:\mathbb{R}^{M}\rightarrow\mathbb{R}$ denotes a scalarisation function and $\rho$ denotes a univariate risk functional (Section 2.3.1). Conceptually, an STR problem associates each input $\mathbf{x}\in\mathbb{X}$ with its risk-adjusted scalarised value: $\rho[s(f(\mathbf{x},\cdot))]\in\mathbb{R}$. The solution of an STR problem is then given by the subset of inputs which map to the largest possible risk-adjusted scalarised value. We refer to these solutions as the STR robust solutions of $f$ (with respect to $s_{\boldsymbol{\theta}}$ and $\rho$). Similar with the RTS method (Section 3.1), the STR method works by reframing the robust multi-objective optimisation problem (1) into a collection of STR problems. As before, each scalarised problem corresponds to one specific preference $\boldsymbol{\theta}\in\Theta$ and one set of solutions that are STR robust with respect to this preference. By solving a collection of these problems, one hopes to identify a collection of STR robust solutions that are desirable for the decision maker. Moreover, as these problems are all scalar-valued, one can in theory appeal to many standard techniques from single-objective optimisation in order to solve them.